1. (-7 square 3)(-4 square 10)
2. square 16/36
3. y/square 7+square 3
Can anyone help with these problems?

Answer by Alan3354(69443)

1. (-7 square 3)(-4 square 10)
= (-7*9)*(-4*100) = -63*-400
= 25200
------------
2. square 16/36
= 256/36
= 64/9
-------
3. y/square 7+square 3
What's the denominator? 1 term or both?
=============
#1 is not clear. Is it -(7)^2 ? or (-7)^2 ? Or 7 times 3^2 ?